Need help with this linear equation word problem.
Use system of equations to solve problem.
A farmer raises wheat and corn on 130 acres of land. He wants to plant 62 more acres of wheat than corn. How many acres of each should he plant.

Answer by stanbon(75887)

A farmer raises wheat and corn on 130 acres of land. He wants to plant 62 more acres of wheat than corn. How many acres of each should he plant.
-------
Acres Equation: w = c + 62
Acres Equation: w + c = 130
----
Substitute for "w" and solve for "c":
c+62 +c = 130
2c = 68
---
c = 34 (# of acres of corn)
---
Solve for "w":
w = c+62
w = 34+62
w = 96 (# of acres of wheat)
================================
Cheers,
Stan H.
------